OpenStudy (anonymous):

Formulas and Identities #1 Question 33 It is given that \(\frac{A}{k-1}-\frac{B}{k-2} = \frac{1}{(k-1)(k-2)}\), where k ≠ 1 and k ≠ 2 (ai) Show that A(k-2) - B(k-1) =1 (aii) Hence, find the values of A and B (b) Hence, find \(\frac{1}{100 \times 102} + \frac{1}{102 \times 104} + \frac{1}{104 \times 106} + ... + \frac{1}{198 \times 200}\) For (a), I've done that. Result of (aii) is A= -1 and B=-1 How to do (b) ?

OpenStudy (anonymous):

i am not sure both A and B = -1, i will bet that this is a telescoping sum

OpenStudy (anonymous):

yes, it is \[\frac{1}{(k-1)(k-2)}=\frac{1}{k-2}-\frac{1}{k-1}\]

OpenStudy (anonymous):

\[\frac12\left(\frac{1}{100} - \frac{1}{102} + \frac{1}{102} + \ldots +\frac{1}{200}\right)\]

OpenStudy (anonymous):

substitute starting with \(k=101\) and see what you get. the left hand side will be the question your are given, the right hand side will telescope

OpenStudy (anonymous):

@Aerospace.Ekaansh why the \(\frac{1}{2}\)?

OpenStudy (anonymous):

I'm lost. Consider the first term for part (b) \[\frac{1}{100 \times 102}\]and use \(\large \frac{1}{(k-1)(k-2)}=\frac{1}{k-2}-\frac{1}{k-1}\) for k=101 \[ \frac{1}{(101-1)(101-2)}= \frac{1}{(100)(99)} = weird\]

OpenStudy (zarkon):

you have some inconsistencies in your original question. I'm betting you have some typos.

OpenStudy (anonymous):

I'm sorry... There are some mistakes in question (ai). It's now corrected

OpenStudy (anonymous):

Ignoring all of the question and just considering the series, \[\frac{1}{100 \times 102} + \frac{1}{102 \times 104} + \frac{1}{104 \times 106} + ... + \frac{1}{198 \times 200} = \sum \frac{1}{k(k+2)}\]\[ = \sum \frac12\left(\frac1k-\frac1{k+2}\right)\]

OpenStudy (anonymous):

Sorry, but how did you get the last thing? And also, I don't think we can ignore all of the question due to the word 'hence'

OpenStudy (zarkon):

notice that all the numbers in the denominator are even...this tells us that \[\sum \frac{1}{k(k+2)}\] probably isn't the sum we want (unless you are letting k increase by 2's)

OpenStudy (anonymous):

Sorry... I meant this: \(\sum \frac12\left(\frac1k-\frac1{k+2}\right)\) And this: \( \sum \frac{1}{k(k+2)}\) should not be something I want for this question.

OpenStudy (anonymous):

I am sorry. :/

OpenStudy (anonymous):

\[\sum\frac {1}{(100 + 2(k-1))(100+2k)}\]?

OpenStudy (zarkon):

that would work...we should put limits on k

OpenStudy (anonymous):

oh wow my answer was all messed up although the partial fractions was right

OpenStudy (zarkon):

I don't see a direct connection from the first part of the question to the second part.

OpenStudy (anonymous):

\[\frac{1}{2k(2k+2)}=\frac{1}{4}\frac{1}{k(k+1)} \] maybe?

OpenStudy (zarkon):

ah

OpenStudy (anonymous):

then use partial fractions, then start at 50 that is my guess but i messed up earlier so who knows

OpenStudy (anonymous):

\[\frac{1}{4}\left(\frac{1}{k}-\frac{1}{k+1}\right)\]

OpenStudy (anonymous):

Thanks I got it

OpenStudy (anonymous):

sorry about messing up the first one good that zarkon is here to keep us honest

OpenStudy (zarkon):

you found the last connection :)

OpenStudy (anonymous):

Hmm.. to be honest.. I got it because I've found my work when you're posting the solution.. Sorry for that... And thanks @Zarkon and @satellite73

OpenStudy (zarkon):

so we are worthless ...man... ;)